OpenStudy (anonymous):

Can you please help me! Find the solution, given the replacement set. 5x - 2y = 6 {(-2, -8), (0, -3), (1, -1), (2, 4), (4, 7)} A. {(-2, -8), (2, 4), (4, 7)} B. {(-2, -8), (0, -3), (4, 7)} C. {(-2, -8), (0, -3), (1, -1)} D. {(1, -1), (2, 4), (4, 7)}

Directrix (directrix):

B. {(-2, -8), (0, -3), (4, 7)} Let's check this one.

Directrix (directrix):

On second thought, perhaps we should check each of the points in the replacement set.

Directrix (directrix):

5x - 2y = 6 {(-2, -8) 5(-2) - 2(-8) = 6 ? -10 + 16 = 6 Yes. (-2, -8) is in the solution set

Directrix (directrix):

Will you test (0, -3) while I check out (1, -1)? Will you?

Directrix (directrix):

(1,-1) 5(1) -2(-1) = 6 ?? 5 + 2 = 6 7=6 NOT in solution set.

Directrix (directrix):

Now, you do ( (2, 4) and I will do (4,7). You can use the drawing tool if you like.

Directrix (directrix):

(4,7) 5(4) - 2(7) = 6? 20 - 14 = 6 6=6 YES (4,7) is in the solution set.

Directrix (directrix):

And, you got what for your points?